# Bounds on detection of Bell correlations with entangled ultra-cold atoms in optical lattices under occupation defects

### v1

@@ V.1 Effective microscopic description
To generate entanglement from the initial spin coherent states (52) driven by the $t$$J$ Hamiltonian (45) we consider two methods. Both of them were thoroughly investigated in Hernández Yanes et al. (2024).
The first (A) uses a weak anisotropy $\Delta\neq 1$ in (46), such that $\Delta\ll 3-2\cos(\pi/M)$ and $\Delta\gg 2\cos(\pi/M)-1$, to generate entanglement in the system. Under an ideal scenario with $N=M$,the system can be represented by a single spin chain whose dynamics is described by the OAT model Hernández Yanes et al. (2024). The presence of holes divides the spin chain into partial chains separated by holes whose configurations are included in (52). For instance, if a single hole is located somewhere in the middle of the chain (not at the borders), the system can be effectively viewed as two partial chains. The system dynamics composed of the partial chains separate only when the positions of holes are fixed. The microscopic model describing the dynamics of each specific configuration involved in (50) is effectively approximated by the OAT model when the positions of holes are pinned to the lattice. The model describing each of $n$ partial chains for a given configuration of $\sigma$ holes reads
@@
where $L_{n}$ is the number of atoms consisting of the partial chain.
@@
with $\beta_{j}\ll J\left[\cos(\frac{\pi}{N})-1\right]$, for the isotropic case $\Delta=1$ to generate entanglement in the system. In this case, when $\sigma$ holes are pinned into the lattice, the effective model describing each of the $n$ partial chains is described effectively by the following Hamiltonian
@@
The effective models for the (A) and (B) methods are derived and explained in detail in Hernández Yanes et al. (2024). Their validity for describing spin-squeezing generation was also demonstrated.
Therefore, one can employ the effective models based on the OAT dynamics for partial chains when evaluating expectation values as given by (54) when considering fixed positions of holes (effectively $t=0$ in the $t$$J$ model (45)). We employ them to calculate the upper bound on the filling factor required for the Bell correlations detection as demonstrated in V.3. However, the effective models become inaccurate when the tunnelling of holes occurs. In such a case, we perform full numerical many-body calculations by the exact diagonalization method of the $t$$J$ model (45) when demonstrating Bell correlations and validity of the Toy model.
In the next subsections, we demonstrate the generation of Bell correlations in the system using scenarios A and B. In our numerical simulations, we consider open boundary conditions Hernández Yanes et al. (2024) and use the parameters as in the recent experiment of A. Rubio-Abadal et al Rubio-Abadal et al. (2019) with 87Rb atoms, lattice spacing $d=532$ nm, and inter- and intraspecies interactions $U_{aa}\sim U_{bb}\sim U_{ab}=U$ Fukuhara et al. (2013) with $U=24.4t$.
@@ V.2 $M$-site Bell correlations
In the upper panel of Fig. 3 we show the evolution of the $M$-site Bell correlator $\mathcal{E}_{M}$ with $M=12$ and for the two proposed methods, A and B. The small shift in time scale and lower magnitude of the correlator maxima in the anisotropic case result from its effective model being approximated to the first order Hernández Yanes et al. (2024). The only relevant contribution comes from the part of the system state describing all sites occupied, as already discussed in Section III. The tunnelling of holes cannot change the value of this correlator.
In the lower panel of Fig. 3 we check the scaling of the correlator with the filling factor $f$ for relevant instances in time using a logarithmic scale. We see good agreement with the scaling of the toy model result in Eq. (16) when identifying $p$ with $f$. Likewise, we see the GHZ state result crossing the classical limit at the value obtained in Eq. (17) for $q=2$, identified by the edge of the grey area.
Figure 3: $M$-sites Bell correlator $\mathcal{E}_{M}=\left|\expectationvalue*{\otimes_{j}^{M}s_{j}^{+}}\right|^{2}$ for the Hamiltonian (45) with $M=12$ sites under scenario A (Anisotropy) with $\Delta=0.98,\beta_{j}=0$ (blue line) and scenario B (Magnetic Field) with $\Delta=1,\beta_{j}=J/50\cos\left[\pi(M-1)/M(j-1/2)\right]$ (orange lines). Upper panel shows the evolution of the correlator for $f=1$. Lower panel shows the scaling at a particular moment in time with the filling factor $f$ for scenario B. The solid line corresponds to the GHZ state, the dashed line to a superposition of $q=4$ coherent states, and the dotted line to the best spin-squeezing time $t_{\rm best}$. The logarithmic scale has been used to illustrate the power-law influence of $f$ and the classical limit $2^{-M}$.
@@ V.3 Two-sites Bell correlations
In Fig. 4 we demonstrate the variation of the two-sites Bell correlations generated using the method A (interactions anisotropy, left column) and B (inhomogeneous magnetic field, right column) for different filling factors and effective tunnelling rates $t/J$. The color bar indicates the magnitude of effective tunnelling when holes are present in the system relative to the relevant energy scale $J$. While in all instances $t\neq 0$ to have a non-zero value of $J$, we assume for the effective tunnelling $t/J=0$ that positions of spins are fixed for those particular numerical simulations. Shaded areas indicate the upper and lower boundaries set by the results of holes fixed in place and holes moving infinitely fast, respectively. The upper bound is calculated numerically for the case of the fixed position of holes by using effective models as described in Hernández Yanes et al. (2024). The lower bound is given by the toy model and the corresponding Bell inequality (III.2.1) where expectation values $\langle{\bf\cdot}\rangle_{\rm SS}$ are replaced with the one given by the corresponding OAT model. We also observe the generation of Bell correlations requires a larger filling factor in the inhomogeneous magnetic field scenario; both when the movement of holes is allowed and when not. Nevertheless, the numerical results demonstrate the effectiveness of the proposed toy model (III.2.1) in the estimation of the lower bound in the detection of two-sites Bell correlations.
@@
Note, the microscopic effective models (55) and (57) differ by the presence of a linear term, such that their results will be bounded by different limits as illustrated in Fig. 5. The critical value for the lower bound in this figure is obtained by solving the corresponding cubic equation by setting (III.2.1) to zero, for which only one real root exists. The anisotropy scenario A, while less accurately described by the corresponding effective OAT model (55) as explained in Hernández Yanes et al. (2024), achieves lower bound results predicted by the toy model (III.2.1) when tunnelling processes happen. This is much harder to achieve in the inhomogeneous magnetic field scenario B since the linear term in (57) changes at each configuration of holes and makes the dynamics largely incoherent. This is a stark contrast with the M-site Bell correlator presented in Fig. 3 where we find the lower bound set by the toy model (16) to be equally accurate.
Figure 5: Critical value of filling factor $f_{c}$ with respect to system size $M$ for the Hamiltonian (45) under scenario A (Anisotropy) with $\Delta=0.98,\beta_{j}=0$ (triangular markers) and scenario B (Mag. Field) with $\Delta=1,\beta_{j}=J/50\cos\left[\pi(M-1)/M(j-1/2)\right]$ (cross markers). The lower and upper bounds are shown in solid lines, with the area between them filled in grey. We use a first passage algorithm with similar data to the one used in Fig. 4 to obtain these values. Due to numerical calculation costs, results for $M=12$ include configurations of all possible numbers of holes in the initial state, while results for $M=14,16$ include configurations with, at least, up to three holes.
@@ VI Entanglement generation in the superfluid phase
We consider here the protocol where the quantum correlations are generated using atom-atom interactions within the superfluid phase and then stored in the Mott phase through adiabatic increasing of the lattice height Kajtoch et al. (2018b); Płodzień et al. (2020b). In this protocol, initially, all atoms occupy the internal state $a$, and at zero temperature, they are in the ground state of the Bose-Hubbard Hamiltonian $|\psi^{(a)}_{0}\rangle$ in the superfluid regime. Subsequently, a $\pi/2$ pulse is applied to put the atoms in a coherent spin state where the collective spin is along the $x$-axis. During the dynamics, the lattice height $V_{0}$ linearly increases as
@@ VI.1 $M$-site Bell correlations
where $P_{0}=e^{-E_{0}^{(0)}/k_{\rm B}T}/Z$ represents the probability of occupying the ground state in (62) [^5] It is important to note that in general, the final external configuration of the system does not consist of a factorized state over lattice sites, as in (8), in particular the probability $P_{0}$ is not of the form $P_{0}=p^{M}$.. Fig. 6-(a) illustrates this relationship for $N=M=4$ over varying temperatures, revealing the existence of a critical temperature $T_{\rm c}$ below which $M$-site Bell correlations are detectable. In the weakly interacting regime, the probability $P_{0}$ can be analytically determined using the Bogoliuobov theory, where, in the initial superfluid regime, when all the atoms occupy the internal state $a$, the system can be approximately described by the Hamiltonian
@@
The critical temperature $T_{\rm c}$, below which Bell correlations in states generated at times $\chi\tau=\pi/q$ of OAT dynamics are detectable, can be determined by setting $P_{0}$ (68) equal to the critical probability $p_{\rm c}^{M}$, where $p_{\rm c}$ is given by (17). Fig. 6-(b), illustrates this critical temperature for a GHZ state (i.e. $q=2$) as a function of the atom number $N$.
Figure 6: Upper panel: $M$-sites Bell correlator $\mathcal{E}_{M}^{(T)}$ at zero and nonzero temperature as a function of the ramp duration $\tau_{\mathrm{ramp}}$ for $N=M=4$ with initial lattice height $V_{\rm ini}/E_{\rm r}=3$ and final lattice height $V_{\rm fin}/E_{\rm r}=20$. Parameters used here are: $U_{aa}=U_{bb}=100.4a_{\rm B}$ and $U_{ab}=0.95U_{aa}$, with $a_{\rm B}$ being the Bohr radius. Colored points represent the exact numerical results, including the $\pi/2$ pulse and the dynamical raising of the lattice, and continuous lines represent Equation (64) where $P_{0}$ was calculated using exact numerical simulations. The horizontal grey dashed line represents the non-locality bound. Lower panel: Critical temperature $T_{\rm c}$, below which the non-local Bell correlations associated to the GHZ state can be detected, as a function of atom number $N$.
@@ VI.2 Two-sites Bell correlations
In large systems, studying $M$-site Bell correlations can be challenging both theoretically and experimentally. Therefore, for such systems, we focus on two-site Bell correlations. In this context, we assume that the system is brought to the Mott phase through a transformation that leaves the system in thermal equilibrium at each moment and conserves the entropy. Under such conditions, one can explicitly write an approximate external density matrix $\hat{\rho}_{\rm ext}$ of the system in the final Mott phase and thus analytically determine the probability $p(T_{\rm i})$ of having exactly one atom in a given lattice site as a function of the initial temperature $T_{\rm i}$. We first calculate the initial entropy $S_{\rm SF}$ as a function of $T_{\rm i}$. We then determine the final temperature $T_{\rm f}$ of the system in Mott using entropy conservation Blakie and Porto (2004); Rey et al. (2006); Yoshimura et al. (2008)
@@
We now calculate the entropy of the system in the Mott phase by considering the limit $t=0$ and small temperatures. In this case, the partition function, when $N=M$, can be estimated using a two particles-holes excitation approximation, where, we take into account only the states with at most two holes and two doubly occupied lattice sites, with $U=U_{\rm aa}$
@@
In numerical simulations of the exact dynamics, $p$ can be calculated as the trace of the projection of the density matrix of the system on the eigenspace of the observable $\hat{n}_{i}$ associated with the eigenvalue $n_{i}=1$. Equation (VI.2) is represented in Fig. 7-(a) as a function of the initial temperature, compared to exact results using both one- and two-component Bose-Hubbard Hamiltonian. This reveals that at sufficiently low temperatures in small systems, the probability $p(T_{\rm i})$ can be accurately approximated using (VI.2) and the entropy conservation condition (69). By equating $p(T_{\rm i})$ with the critical probability (41) from the toy model, one can determine an upper bound, dependent on the ramp duration $\tau_{\mathrm{ramp}}$, for the initial temperature $T_{\rm c}$ above which two-sites non-local Bell correlations cannot be detected. Fig. 7-(b) shows the critical temperature associated to the minimal (critical) $p$, over the ramp duration $\tau_{\mathrm{ramp}}$, as a function of the atom number $N$   [^6] For larger atoms number, beyond the range of $N$ presented in the figure, the two-hole approximation becomes invalid..

### v2

@@ V.1 Effective microscopic description
To generate entanglement from the initial spin coherent states (52) driven by the $t$$J$ Hamiltonian (45) we consider two methods. Both of them were thoroughly investigated in [41].
The first (A) uses a weak anisotropy $\Delta\neq 1$ in (46), such that $\Delta\ll 3-2\cos(\pi/M)$ and $\Delta\gg 2\cos(\pi/M)-1$, to generate entanglement in the system. Under an ideal scenario with $N=M$, the system can be represented by a single spin chain whose dynamics is described by the OAT model [41]. The presence of holes divides the spin chain into partial chains separated by holes whose configurations are included in (52). For instance, if a single hole is located somewhere in the middle of the chain (not at the borders), the system can be effectively viewed as two partial chains. The system dynamics composed of the partial chains separate only when the positions of holes are fixed. The microscopic model describing the dynamics of each specific configuration involved in (50) is effectively approximated by the OAT model when the positions of holes are pinned to the lattice. The model describing each of $n$ partial chains for a given configuration of $\sigma$ holes reads
@@
where $L_{n}$ is the number of atoms composing the partial chain.
@@
with $\beta_{j}\ll J\left[\cos(\frac{\pi}{M})-1\right]$, for the isotropic case, $\Delta=1$ to generate entanglement in the system. In this case, when $\sigma$ holes are pinned into the lattice, the effective model describing each of the $n$ partial chains is described effectively by the following Hamiltonian
@@
The effective models for the (A) and (B) methods are derived and explained in detail in [41]. Their validity for describing spin-squeezing generation was also demonstrated.
Therefore, one can employ the effective models based on the OAT dynamics for partial chains for evaluating expectation values as given by (54) when considering fixed positions of holes (effectively $t=0$ in the $t$$J$ model (45)). As we show it in V.3, this model provides an upper bound on the filling factor required for the Bell correlations detection. The model, however, becomes inaccurate when the tunnelling of holes occurs. In such a case, to test the validity of the toy model, we perform full numerical many-body calculations by the exact diagonalization method of the $t$$J$ model (45).
In the next subsections, we demonstrate the generation of Bell correlations in the system using scenarios A and B. In our numerical simulations, we consider open boundary conditions [41] and use the parameters as in the recent experiment of A. Rubio-Abadal et al [48] with 87Rb atoms, lattice spacing $d=532$ nm, and inter- and intraspecies interactions $U_{aa}\sim U_{bb}\sim U_{ab}=U$ [53] with $U=24.4t$.
@@ V.2 $M$-site Bell correlations
In the upper panel of Fig. 3 we show the evolution of the $M$-site Bell correlator $\mathcal{E}_{M}$ with $M=12$ and for the two proposed methods, A and B. For the $M$-site correlator the only relevant contribution in (52) comes from the part of the system state describing all sites occupied, as already discussed in Section III. The tunnelling of holes cannot change the value of this correlator.
In the lower panel of Fig. 3 we check the scaling of the correlator with the filling factor $f$ for relevant instants in time, $\chi\tau=\pi/2$ (solid line) corresponding to the GHZ state. We see good agreement with the scaling of the toy model result in Eq. (16) when identifying $p$ with $f$. Likewise, we see the GHZ state result crossing the classical limit at the value obtained in Eq. (17) for $q=2$, identified by the edge of the grey area.
Figure 3: $M$-sites Bell correlator for the Hamiltonian (45) with $M=12$ sites under scenario A (Anisotropy) with $\Delta=0.98,\beta_{j}=0$ (blue line) and scenario B (Magnetic Field) with $\Delta=1,\beta_{j}=J/50\cos\left[\pi(M-1)/M(j-1/2)\right]$ (orange lines). The upper panel shows the evolution of the correlator for $f=1$. The lower panel shows the scaling at three particular moments in time with the filling factor $f$ for scenario B. The solid line corresponds to the GHZ state, the dashed line to a superposition of $q=4$ coherent states, and the dotted line to the best spin-squeezing time $t_{\rm best}$. The logarithmic scale has been used to illustrate the power-law dependence of $f$ and the classical limit $2^{-M}$.
@@ V.3 Two-sites Bell correlations
Figure 5: The critical value of filling factor $f_{c}$ concerning the system size $M$ for the Hamiltonian (45) under scenario A (Anisotropy) with $\Delta=0.98,\beta_{j}=0$ (triangular markers) and scenario B (Mag. Field) with $\Delta=1,\beta_{j}=J/50\cos\left[\pi(M-1)/M(j-1/2)\right]$ (cross markers). The lower and upper bounds are shown in solid lines, with the area between them filled in grey. We numerically find $f_{c}$ from similar data to the ones shown in Fig. 4. Due to numerical calculation costs, results for $M=12$ include configurations of all possible numbers of holes in the initial state, while results for $M=14,16$ include configurations with at least three holes in (50).
In Fig. 4 we demonstrate the variation of the two-sites Bell correlations generated using the method A (interactions anisotropy, left column) and B (inhomogeneous magnetic field, right column) for different filling factors and effective tunnelling rates $t/J$. The color bar indicates the magnitude of effective tunnelling when holes are present in the system relative to the relevant energy scale $J$. While in all instances $t\neq 0$ to have a non-zero value of $J$, we assume for the effective tunnelling $t/J=0$ that positions of spins are fixed for those particular numerical simulations. Shaded areas indicate the upper and lower boundaries set by the results of holes fixed in place and holes moving infinitely fast, respectively. The upper bounds are calculated numerically for the case of the fixed position of holes by using effective models as described in [41]. Note that the microscopic effective models (55) and (57) differ by the presence of a linear term, such that their results will be bounded by different limits as illustrated in Fig. 5. The lower bound is given by the toy model and the corresponding Bell inequality (III.2.1) where expectation values $\langle{\bf\cdot}\rangle_{\rm SS}$ are replaced with the one given by the corresponding OAT model. We observe that the generation of Bell correlations is possible with a lower filling factor in the anisotropy method (A), both when the movement of holes is allowed and when not. Only in this case, the numerical results demonstrate the effectiveness of the proposed toy model (III.2.1) in the estimation of the lower bound in the detection of two-sites Bell correlations. In the case of inhomogeneous magnetic field method (B) we could not reach the lower bound within the explored region of parameters.
The critical value for the lower bound in this figure is obtained by solving the cubic equation obtained by setting (III.2.1) to zero, for which only one real root exists. The anisotropy scenario A, while less accurately described by the corresponding effective OAT model (55) as explained in [41], achieves the lower bound results predicted by the toy model (III.2.1) for a sufficiently large ratio $t/J=U_{ab}/4t$. This is much harder to achieve in the inhomogeneous magnetic field scenario B since the linear term in (57) changes at each configuration of holes and makes the dynamics largely incoherent. This is a stark contrast with the M-site Bell correlator presented in Fig. 3 where we find the lower bound set by the toy model (16) to be accurate for the two scenarios.
@@ VI Entanglement generation in the superfluid phase
We consider here the protocol where the quantum correlations are generated using atom-atom interactions within the superfluid phase and then stored in the Mott phase through adiabatic increasing of the lattice height [43, 54]. We assume that the tunnelling parameter $t$ and interaction parameters $U_{\sigma\sigma^{\prime}}$ can by approximated as [54]
$t$ $\approx\frac{4}{\sqrt{\pi}}\left(\frac{V_{0}}{E_{r}}\right)^{3/4}e^{-2\sqrt{{V_{0}}/{E_{r}}}},$ (61)
$U_{\sigma\sigma^{\prime}}$ $\approx\sqrt{\frac{32}{\pi}}\frac{a_{\sigma\sigma^{\prime}}d}{L_{\perp}^{2}}\left(\frac{V_{0}}{E_{r}}\right)^{1/4},$ (62)
where $a_{\sigma\sigma^{\prime}}$ are the s-wave scattering lengths, $V_{0}$ is the lattice potential depth and $L_{\perp}$ is the characteristic length of the wave function perpendicular to the lattice direction.
In this protocol, initially, all atoms occupy the internal state $a$, and at zero temperature, they are in the ground state of the Bose-Hubbard Hamiltonian $|\psi^{(a)}_{0}\rangle$ in the superfluid regime. Subsequently, a $\pi/2$ pulse is applied to put the atoms in a coherent spin state where the collective spin is along the $x$-axis. During the dynamics, the lattice height $V_{0}$ linearly increases as
@@ VI.1 $M$-site Bell correlations
where $P_{0}=e^{-E_{0}^{(0)}/k_{\rm B}T}/Z$ represents the probability of occupying the ground state in (64) [^5] It is important to note that in general, the final external configuration of the system does not consist of a factorized state over lattice sites, as in (8), in particular the probability $P_{0}$ is not of the form $P_{0}=p^{M}$.. The upper panel of Fig. 6 illustrates this relationship for $N=M=4$ over varying temperatures, revealing the existence of a critical temperature $T_{\rm c}$ below which $M$-site Bell correlations are detectable. In the weakly interacting regime, the probability $P_{0}$ can be analytically determined using the Bogoliuobov theory, where, in the initial superfluid regime, when all the atoms occupy the internal state $a$, the system can be approximately described by the Hamiltonian
@@
The critical temperature $T_{\rm c}$, below which Bell correlations in states generated at times $\chi\tau=\pi/q$ of OAT dynamics are detectable, can be determined by setting $P_{0}$ (70) equal to the critical probability $P_{0}=p_{\rm c}^{M}$, where $p_{\rm c}$ is given by (17). The lower panel of Fig. 6 illustrates this critical temperature for a GHZ state (i.e. $q=2$) as a function of the atom number $N$. One can find analytically the critical temperature for $N\to\infty$ by taking the continuous limit in (70), which yields
[…]
We again take $P_{0}=p_{c}^{M}$ for $q=2$ to obtain the critical temperature
$k_{\mathrm{B}}T_{c}=\frac{3\ln{2}}{\pi}\sqrt{2tU},$ (72)
which is illustrated as a dashed line in the lower panel of Fig. 6.
Figure 6: Upper panel: $M$-sites Bell correlator $\mathcal{E}_{M}^{(T)}$ at zero and nonzero temperature as a function of the ramp duration $\tau_{\mathrm{ramp}}$ for $N=M=4$ with initial lattice height $V_{\rm ini}/E_{\rm r}=3$ and final lattice height $V_{\rm fin}/E_{\rm r}=20$. Parameters used here are: $a_{a}=a_{b}=100.4a_{\rm B}$ and $a_{ab}=0.95a_{a}$, with $a_{\rm B}$ being the Bohr radius, $d=431$ nm, $L_{\perp}=d/\sqrt{2\pi}$. Colored points represent the exact numerical results, including the $\pi/2$ pulse and the dynamical raising of the lattice, and the colored solid lines superimposed to the symbols represent Equation (66) where $P_{0}$ was calculated using exact numerical simulations. The horizontal grey dashed line represents the non-locality bound. Lower panel: Critical temperature $T_{\rm c}$, below which the non-local Bell correlations associated with the GHZ state can be detected, as a function of atom number $N$. The dashed line corresponds to the large $N$ limit (72).
@@ VI.2 Two-sites Bell correlations
Here, we focus on two-site correlations that can be measured also in large systems. In this context, we assume that the system is brought to the Mott phase through a transformation that leaves the system in thermal equilibrium at each moment and conserves the entropy. Under such conditions, one can explicitly write an approximate external density matrix $\hat{\rho}_{\rm ext}$ of the system in the final Mott phase and thus analytically determine the probability $p(T_{\rm i})$ of having exactly one atom in a given lattice site as a function of the initial temperature $T_{\rm i}$. We first calculate the initial entropy $S_{\rm SF}$ as a function of $T_{\rm i}$. We then determine the final temperature $T_{\rm f}$ of the system in Mott using entropy conservation [55, 56, 57]
@@
We now calculate the entropy of the system in the Mott phase by considering the limit $t\to 0$ and small temperatures. In this case, the partition function, when $N=M$, can be estimated using a two particles-holes excitation approximation, where, we take into account only the states with at most two holes and two doubly occupied lattice sites, with $U=U_{aa}$
@@
In numerical simulations of the exact dynamics, $p$ can be calculated as the trace of the projection of the density matrix of the system on the eigenspace of the observable $\hat{n}_{i}$ associated with the eigenvalue $n_{i}=1$. Equation (VI.2) is represented in the upper panel of Fig. 7 as a function of the initial temperature, compared to exact results using both one- and two-component Bose-Hubbard Hamiltonian. This reveals that at sufficiently low temperatures in small systems, the probability $p(T_{\rm i})$ can be accurately approximated using (VI.2) and the entropy conservation condition (73). By equating $p(T_{\rm i})$ with the critical probability (41) from the toy model, one can determine an upper bound, dependent on the ramp duration $\tau_{\mathrm{ramp}}$, for the initial temperature $T_{\rm c}$ above which two-sites non-local Bell correlations cannot be detected. The lower panel of Fig. 7 shows the critical temperature associated to the smallest critical $p$, minimized over the ramp duration $\tau_{\mathrm{ramp}}$, as a function of the atom number $N$   [^6] For larger atoms number, beyond the range of $N$ presented in the figure, the two-hole approximation becomes invalid..